# No Minima, No Collisions: Combining Modulation and Control Barrier Function Strategies for Feasible Dynamical Collision Avoidance

### v1

@@ VIII-B Robot Dynamics
In tests using fully actuated systems, the performances of the proposed reference and on-manifold MCBF-QP methods in 2-dimensional Euclidean space (R-MCBF-QP and onM-MCBF-QP) compete against regular CBF-QP and the proposed constrained Mod-DS using standard Ridgeback dynamics below.
$\begin{bmatrix}\dot{x}\\ \dot{y}\end{bmatrix}=\begin{bmatrix}u_{x}\\ u_{y}\end{bmatrix}$ (74)
In tests for underactuated control affine systems, the performance of the proposed on-manifold MCBF-QP method in 2-dimensional Euclidean space (shifted on-manifold Mod-based CBF-QP, S-onM-MCBF) competes against regular CBF-QP using the shifted unicycle model (shifted CBF-QP, S-CBF-QP), while on-manifold MCBF-QP method’s performance in 3-dimensional non-Euclidean space (high order on-manifold Mod-based CBF-QP, HO-onM-MCBF) competes against higher order CBF-QP (HO-CBF-QP) using Fetch dynamics. The most commonly used model for differential drive robots like Fetch is the unicycle model below.
$\dot{\xi}=\begin{bmatrix}\dot{x}\\ \dot{y}\\ \dot{\theta}\end{bmatrix}=\begin{bmatrix}\cos{\theta}&0\\ \sin{\theta}&0\\ 0&1\end{bmatrix}\begin{bmatrix}v\\ \omega\end{bmatrix}$ (75)
@@
$\dot{\xi}=\begin{bmatrix}\dot{x}\\ \dot{y}\\ \dot{\theta}\end{bmatrix}=\begin{bmatrix}\cos{\theta}&-a\sin{\theta}\\ \sin{\theta}&a\cos{\theta}\\ 0&1\end{bmatrix}\begin{bmatrix}v\\ \omega\end{bmatrix}$ (76)
@@
$\frac{\partial h(x,y,t)}{\partial x}\dot{x}+\frac{\partial h(x,y,t)}{\partial y}\dot{y}+\frac{\partial h(x,y,t)}{\partial t}\geq-\alpha(h(x,y,t)).$ (77)
While the shifted unicycle model enables CBFsto be defined as signed distance functions like $h_{\text{conv}}$, $h_{\text{star}}$ and $h_{\text{nstar}}$, the method cannot be easily generalized to all underactuated robots. HO-CBF-QP solves the generalizability issue by using high order control barrier functions (HO-CBFs) $h_{\text{OH}}(\xi,t)$ of relative degree 1 to control inputs, instead of signed distance functions $h(x,y,t)$ as CBFs. Here, we demonstrate that the on-manifold MCBF-QP method we proposed, likewise can be generalized to high-relative-order robot dynamics by controlling Fetch through the standard unicycle model in (75) using high-order on Manifold Mod-based CBF-QP (HO-onM-MCBF). HO-CBF $h_{\text{OH}}(\xi,t)$ used by 3-dimensional HO-onM-MCBF and HO-CBF-QP in our validations are constructed as the following, where $w\in\mathbb{R}^{+}$.
$h_{\text{OH}}(\xi,t)=h(x,y,t)+w\frac{\partial h(x,y,t)}{\partial x}\cos{\theta}+w\frac{\partial h(x,y,t)}{\partial y}\sin{\theta}$ (78)
@@ VIII-C Geodesic Approximation in Non-Euclidean Space
[16] shows that the obstacle exit strategy of geodesic approximation, given well-tuned step size $\beta$ and horizon size $N$, can generate a first-order approximation to the obstacle surface in Euclidean space $\mathbb{R}^{d}$ alongside an effective circumventing strategy, connecting robot current position to target position with the least reward $P_{N}$. In this work, we demonstrate that such a strategy can be extended to non-Euclidean space, by replacing $x_{i}$ with $\xi_{i}$, when used in Mod-based CBF-QP, making possible local-minimum-free control affine CBF-QP in general.
Here, the application of the geodesic approximation strategy is validated using the 3-dimensional non-Euclidean state space for the unicycle model $[x,y,\theta]$. Figure 11(a), (b) and (c) shows what the unsafe region defined by HO-CBF in Equation 78 looks like in our constructed 3D non-Euclidean space. For visual effects and interpretability of the obstacle, only the section with $\theta\in[0,2\pi)$ is shown. Given robot initial state $\xi_{0}=[3,3,0]$ and 20 $e^{0}$ vectors evenly distributed in the plane tangent to the gradient of obstacle $h_{\text{OH}}(\xi_{0},t)$, geodesic approximation strategy, as shown in Figure 11, is able to select the circumventing trajectories traveling most efficiently towards the target using reward function $p(\xi_{i},x^{*})=||\xi_{i}[:2]-x^{*}||_{2}$.
@@ VIII-E Real World Experiments
In addition to simulations, the proposed S-onM-MCBF-QP and HO-onM-MCBF-QP methods are compared to S-CBF-QP and HO-CBF-QP in real-life experiments using Fetch given nominal controller in (79). In the first experiment, Fetch is challenged to reach the goal position on the other side of the room despite active attempts from a walking human to block its path. In the second experiment, the robot is tasked to navigate out of a C-shaped cluster region towards the target. All experiments are repeated 5 times with different robot initial poses in static scenarios and different human walking trajectories in dynamic scenarios. The results demonstrate that CBF-induced local minima will deteriorate robot task completion performance, while the introduction of onM-Mod-based constraints into the same QP problem can sufficiently eliminate those local minima. In 2D navigation tasks using Fetch, though both HO-onM-MCBF-QP and S-onM-MCBF-QP drastically improve robot task completion performance, S-onM-MCBF-QP using the shifted unicycle model outperforms HO-onM-MCBF-QP using the standard unicycle model in terms of path efficiency and task completion rate. The reason for that is the unsafe region defined by HOCBF $h_{\text{OH}}$ (78), as shown in 11(c), is not the same as the actual obstacle region, causing the geodesic approximation strategy to output sometimes non-optimal choices of tangent vectors $\phi(\xi,t)$ for the robot to track. Additionally, finding appropriate step size $\beta$ and horizon $N$ to use during geodesic approximation are more challenging for HO-MCBF-QP that searches for $\phi(\xi,t)$ in 3D non-Euclidean state spaces, compared with S-MCBF-QP that looks for $\phi(x,y,t)$ instead in the 2D Euclidean state spaces. However, HO-onM-MCBF-QP’s merit is its generalizability which can be theoretically adopted by all robot models.

### v2

@@ VIII-B Robot Dynamics
In tests using fully actuated systems, the performances of the proposed reference and on-manifold MCBF-QP methods in 2-dimensional Euclidean space (R-MCBF-QP and onM-MCBF-QP) compete against regular CBF-QP and the proposed constrained Mod-DS using standard Ridgeback dynamics below, where $p_{x}$ and $p_{y}$ are robot locations in $x$ and $y$ axis.
$\begin{bmatrix}\dot{p}_{x}\\ \dot{p}_{y}\end{bmatrix}=\begin{bmatrix}u_{x}\\ u_{y}\end{bmatrix}$ (78)
In tests for underactuated control affine systems, the performance of the proposed on-manifold MCBF-QP method in 2-dimensional Euclidean space (shifted on-manifold Mod-based CBF-QP, S-onM-MCBF) competes against regular CBF-QP using the shifted unicycle model (shifted CBF-QP, S-CBF-QP), while on-manifold MCBF-QP method’s performance in 3-dimensional non-Euclidean space (high order on-manifold Mod-based CBF-QP, HO-onM-MCBF) competes against higher order CBF-QP (HO-CBF-QP) using Fetch dynamics. The most commonly used model for differential drive robots like Fetch is the unicycle model below, where $\theta$ is robot orientation in radians measured with respect to the positive $x$ axis.
$\dot{x}=\begin{bmatrix}\dot{p_{x}}\\ \dot{p}_{y}\\ \dot{\theta}\end{bmatrix}=\begin{bmatrix}\cos{\theta}&0\\ \sin{\theta}&0\\ 0&1\end{bmatrix}\begin{bmatrix}v\\ \omega\end{bmatrix}$ (79)
@@
$\dot{x}=\begin{bmatrix}\dot{p}_{x}\\ \dot{p}_{y}\\ \dot{\theta}\end{bmatrix}=\begin{bmatrix}\cos{\theta}&-a\sin{\theta}\\ \sin{\theta}&a\cos{\theta}\\ 0&1\end{bmatrix}\begin{bmatrix}v\\ \omega\end{bmatrix}$ (80)
@@
$\frac{\partial h(p_{x},p_{y},x_{o})}{\partial p_{x}}\dot{p}_{x}+\frac{\partial h(p_{x},p_{y},x_{o})}{\partial y}\dot{p}_{y}+\nabla_{x_{o}}h(x,x_{o})\dot{x}_{o}\geq$ (81)
$-\alpha(h(p_{x},p_{y},x_{o})).$
While the shifted unicycle model enables CBFs to be defined as signed distance functions like $h_{\text{conv}}$, $h_{\text{star}}$ and $h_{\text{nstar}}$, the method cannot be easily generalized to all underactuated robots. HO-CBF-QP solves the generalizability issue by using high order control barrier functions (HO-CBFs) $h_{\text{OH}}(x,x_{o})$ of relative degree 1 to control inputs, instead of signed distance functions $h(p_{x},p_{y},x_{o})$ as CBFs. Here, we demonstrate that the on-manifold MCBF-QP method we proposed, likewise can be generalized to high-relative-order robot dynamics by controlling Fetch through the standard unicycle model in (79) using high-order on Manifold Mod-based CBF-QP (HO-onM-MCBF). HO-CBF $h_{\text{OH}}(x,t)$ used by 3-dimensional HO-onM-MCBF and HO-CBF-QP in our validations are constructed as the following, where $w\in\mathbb{R}^{+}$.
$h_{\text{OH}}(x,x_{o})=h(p_{x},p_{y},x_{o})+w\frac{\partial h(p_{x},p_{y},x_{o})}{\partial p_{x}}\cos{\theta}$ (82)
$+w\frac{\partial h(p_{x},p_{y},x_{o})}{\partial p_{y}}\sin{\theta}$
HO-CBF constraints in CBF-QP and MCBF-QP can be obtained by substituting $h_{OH}(x,x_{o})$ as $h(x,x_{o})$ directly into CBF constraint inequality in Eq (75).
@@ VIII-C Geodesic Approximation in Non-Euclidean Space
[16] shows that the obstacle exit strategy of geodesic approximation, given well-tuned step size $\beta$ and horizon size $N$, can generate a first-order approximation to the obstacle surface in Euclidean space $\mathbb{R}^{d}$ alongside an effective circumventing strategy, connecting robot current position to target position with the least reward $P_{N}$. In this work, we demonstrate that such a strategy can be extended to state $x$ in non-Euclidean space when used in Modulated CBF-QP, making possible local-minimum-free control affine CBF-QP in general.
Here, the application of the geodesic approximation strategy is validated using the 3-dimensional non-Euclidean state space for the unicycle model $[p_{x},p_{y},\theta]$. Figure 11(a), (b) and (c) shows what the unsafe region defined by HO-CBF in Equation 82 looks like in our constructed 3D non-Euclidean space. For visual effects and interpretability of the obstacle, only the section with $\theta\in[0,2\pi)$ is shown. Given robot initial state $x^{0}=[3,3,0]$ and 20 $e^{0}$ vectors evenly distributed in the plane tangent to the gradient of obstacle $h_{\text{OH}}(x^{0},x_{o})$, geodesic approximation strategy, as shown in Figure 11, is able to select the circumventing trajectories traveling most efficiently towards the target using reward function $p(x^{i},x^{*})=||x^{i}[:2]-x^{*}||_{2}$.
@@ VIII-E Real World Experiments
In addition to simulations, the proposed S-onM-MCBF-QP and HO-onM-MCBF-QP methods are compared to S-CBF-QP and HO-CBF-QP in real-life experiments using Fetch given nominal controller in (83). In the first experiment, Fetch is challenged to reach the goal position on the other side of the room despite active attempts from a walking human to block its path. In the second experiment, the robot is tasked to navigate out of a C-shaped cluster region towards the target. All experiments are repeated 5 times with different robot initial poses in static scenarios and different human walking trajectories in dynamic scenarios. The results demonstrate that CBF-induced local minima will deteriorate robot task completion performance, while the introduction of onM-Mod-based constraints into the same QP problem can sufficiently eliminate those local minima. In 2D navigation tasks using Fetch, though both HO-onM-MCBF-QP and S-onM-MCBF-QP drastically improve robot task completion performance, S-onM-MCBF-QP using the shifted unicycle model outperforms HO-onM-MCBF-QP using the standard unicycle model in terms of path efficiency and task completion rate. The reason for that is the unsafe region defined by HOCBF $h_{\text{OH}}(x,x_{o})$ (82), as shown in 11(c), is not the same as the actual obstacle region, causing the geodesic approximation strategy to output sometimes non-optimal choices of tangent vectors $\phi(x,x_{o})$ for the robot to track. Additionally, finding appropriate step size $\beta$ and horizon $N$ to use during geodesic approximation are more challenging for HO-MCBF-QP that searches for $\phi(x,x_{o})$ in 3D non-Euclidean state spaces, compared with S-MCBF-QP that looks for $\phi(p_{x},p_{y},x_{o})$ instead in the 2D Euclidean state spaces. However, HO-onM-MCBF-QP’s merit is its generalizability which can be theoretically adopted by all robot models.